OpenStudy (anonymous):

Find the missing factor............................ x^8=(x^5)(________)

OpenStudy (anonymous):

The first part of this should help out where it shows simplification :) https://www.purplemath.com/modules/exponent.htm

OpenStudy (anonymous):

That did nothing.. aha.. i still dont get it.

OpenStudy (anonymous):

Well... you can treat the exponential terms as addition, so think of the problem like this: x^8=(x^5)(x^y) Now with exponential addition rules, you can say 8=5+y Solve for y and that is your missing term :)